according to study the average number of hours of tv of viewing per household per week is philippines is 50.4 hours and the standard deviation is 11.8 hours.if a random sample of 42 households is taken what is the probability that the sample average is less than 47.5 hours

1
Expert's answer
2022-03-31T05:51:11-0400

"P(\bar X<47.5)=P(Z<\frac{47.5-50.4}{\frac{11.8}{\sqrt{42}}})=P(Z<-1.59)=0.0559."